OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). cos x = sin 2x

OpenStudy (anonymous):

is this \[\cos(x)=\sin(2x)\] or is it \[\cos(x)=\sin^2(x)\]?

OpenStudy (anonymous):

The first one!

OpenStudy (anonymous):

use the identity \[\sin(2x)=2\sin(x)\cos(x)\] to rewrite it as \[\cos(x)=2\sin(x)\cos(x)\]

OpenStudy (anonymous):

or better yet \[2\sin(x)\cos(x)-\cos(x)=0\] or still better \[\cos(x)\left(2\sin(x)-1\right)=0\]then it will be much easier

OpenStudy (anonymous):

How on Earth did you do that? How do cos(x) and sin(2x) equal the same thing?

OpenStudy (anonymous):

step one is an identity \[\sin(2x)=2\sin(x)\cos(x)\] always

OpenStudy (anonymous):

step to is taking \[\cos(x)=2\sin(x)\cos(x)\] and set it equal to zero

OpenStudy (anonymous):

Wouldn't it go to 2sinx=0?

OpenStudy (anonymous):

subtract \(\cos(x)\) from both sides and get \[2\sin(x)\cos(x)-\cos(x)=0\]

OpenStudy (anonymous):

Oh, my bad. Confusion between the multiplication and addition. What then?

OpenStudy (anonymous):

How do you find solutions for one like this? Factor?